OpenStudy (ricktoraw):

Solve for x. 5x + 17 = 3x + 7 A. x = 10 B. x = 5 C. x = –10 D. x = –5

OpenStudy (ricktoraw):

Solve for y. 3y + 7 = y + 13 A. y = 3 B. y = 4 C. y = 5 D. y = 12

OpenStudy (ricktoraw):

Solve for x. 4(2x + 1) = 3(x + 8) A. x = 20 B. x = 7 C. x = 5 D. x = 4

OpenStudy (ricktoraw):

olve for m. 2(5m + 4) = 2(3m – 10) A. m = –28 B. m = –7 C. m = –14 D. m = 14

OpenStudy (ricktoraw):

olve for y. 5(y + 7) = 10y A. y = 7 B. y = –7 C. y = 35 D. y = –35

OpenStudy (welshfella):

Hi the first question solve for x 5x + 17 = 3x + 7 we need to bring all the terms in x to left side of the equation and the numbers on the right so first subtract 3x from both sides ( we have to do same operation both sides):- 5x - 3x + 17 = 3x - 3x + 7 Simplifying will remove 3x from left side:- 2x + 17 = 7 OK soo far?

OpenStudy (welshfella):

* remove 3x from right side

OpenStudy (welshfella):

do you follow the above ok?

OpenStudy (ricktoraw):

no i dont just need answer i will meadal and add u as freiend

OpenStudy (ricktoraw):

i need answer

OpenStudy (welshfella):

I cant give you an answer

OpenStudy (welshfella):

its against Code of Conduct

OpenStudy (ricktoraw):

help

OpenStudy (welshfella):

2x + 17 = 7 subtract 17 from both sides 2x + 17 - 17 = 7 - 17 2x = -10 now we have only terms in x on one side and a number on right to solve for x you now divide both sides by 2 so x = -10 / 2

OpenStudy (devonhoward15):

plus theres like 4 questions and people want a medal for each question anwsered

OpenStudy (ricktoraw):

ok